# pAGN: the one-stop solution for AGN disc modeling

### v1

@@ 4.2 Migration traps
Figure 8 shows the migration-torque profiles for a $M=10^{6}M_{\odot}$ SMBH in both the Sirko & Goodman (2003) and Thompson et al. (2005) model. We use $\epsilon_{\mathrm{S}}=0.1$, $\alpha=0.1$, $l_{\mathrm{E}}=0.5$ and $b=0$ for the SG03 AGN disc and $r_{\rm out}=10^{7}R_{\mathrm{s}}$, $\epsilon_{\mathrm{T}}=10^{-3}$, $\xi=1$, $m_{\rm T}=0.2$, and $\dot{M}_{\rm out}=1.5\times 10^{-2}\,M_{\odot}\,{\rm yr}^{-1}$ for the TQM05 model. The outer accretion rate $\dot{M}_{\rm out}$ was set smaller than the value given by Eq. (27) in order to enforce $h/r<1$ throughtout the disc, unlike the small mass case in Fig. 5. Identifying a migration trap corresponds to regions of the disc where the net migration torque is zero and goes from negative (i.e. outward migration) to positive (i.e. inward migration) as $r$ increases.
The top panel shows the migration torque using both Eq. (29) by Paardekooper et al. (2010) and Eq. (30) by Jiménez & Masset (2017). When using the former, we find migration traps at $r\approx 5\times 10R_{\rm s}$ and $r\approx 10^{3}R_{\rm s}$ for the SG03 model, which is in line with the results reported by both Bellovary et al. (2016) and Grishin et al. (2023). When using the updated migration torque values by Jiménez & Masset (2017) for the SG03 model, we find that the migration torque is always negative and thus the migrator moves across the disc without being trapped. This result is in agreement with those by Grishin et al. (2023). Once the thermal torque from Eq. (34) is added to the updated migration torque of Eq. (30), the bottom panel in Fig. 8 shows that we again obtain migration traps. In the SG03 AGN disc, we find a migration trap for a $M=10^{6}\,{\rm M}_{\odot}$ central BH and a $10M_{\odot}$ migrator occurring at $r\approx 4.77\times 10^{2}R_{\rm s}=3.27\times 10^{-5}$ pc.
When considering the Thompson et al. (2005) model, we obtain a larger number of migration traps, irrespective of the torque prescriptions adopted and the inclusion of the thermal torque contribution to the net torque. For the migration torque by Jiménez & Masset (2017) (the top panel in Fig. 8), we find migration traps form in the TQM05 disc when the gradient $\mathrm{d}\ln\Sigma_{\rm g}/\mathrm{d}\ln r$ discretely changes values, as can be seen in the lower panels of Fig. 8 at $r\approx 7.5\times 10^{2}R_{\rm s}$ and $r\approx 10^{4}R_{\rm s}$. When both migration and thermal torques are considered, we find traps at $r\approx 3.5\times 10^{3}R_{\rm s}$, $r\approx 10^{4}R_{\rm s}$ and $r\approx 3.5\times 10^{5}R_{\rm s}$ for the Thompson et al. (2005) model.
@@ Acknowledgements
We thank Maria Paola Vaccaro, Sean McGee, and Cressida Cleland for discussions. D.Gan. and D.Ger. are supported by ERC Starting Grant No. 945155–GWmining, Cariplo Foundation Grant No. 2021-0555, MUR PRIN Grant No. 2022-Z9X4XS, and the ICSC National Research Centre funded by NextGenerationEU. A.A.T. acknowledges support from JSPS KAKENHI Grant No. 21K13914 and from the European Union’s Horizon 2020 and Horizon Europe research and innovation programs under the Marie Skłodowska-Curie grant agreements No. 847523 and No. 101103134. D.Ger. is supported by MSCA Fellowships No. 101064542–StochRewind and No. 101149270–ProtoBH. Computational work was performed at CINECA with allocations through INFN and Bicocca.

### v2

@@ 4.2 Migration traps
Figure 8 shows the migration-torque profiles for a $M=10^{6}M_{\odot}$ SMBH in both the Sirko & Goodman (2003) and Thompson et al. (2005) model. We use $\epsilon_{\mathrm{S}}=0.1$, $\alpha=0.01$, $l_{\mathrm{E}}=0.5$ and $b=0$ for the SG03 AGN disc and $r_{\rm out}=10^{7}R_{\mathrm{s}}$, $\epsilon_{\mathrm{T}}=10^{-3}$, $\xi=1$, $m_{\rm T}=0.2$, and $\dot{M}_{\rm out}=1.5\times 10^{-2}\,M_{\odot}\,{\rm yr}^{-1}$ for the TQM05 model. The outer accretion rate $\dot{M}_{\rm out}$ was set smaller than the value given by Eq. (27) in order to enforce $h/r<1$ throughtout the disc, unlike the small mass case in Fig. 5. Identifying a migration trap corresponds to regions of the disc where the net migration torque is zero and goes from negative (i.e. inward migration) to positive (i.e. outward migration) as $r$ increases.
The top panel shows the migration torque using both Eq. (29) by Paardekooper et al. (2010) and Eq. (30) by Jiménez & Masset (2017). When using the former, we find migration traps at $r\approx 22R_{\rm s}$ and $r\approx 10^{3}R_{\rm s}$ for the SG03 model, which is in line with the results reported by both Bellovary et al. (2016) and Grishin et al. (2024). When using the updated migration torque values by Jiménez & Masset (2017) for the SG03 model, we find that the migration torque is always negative and thus the migrator moves across the disc without being trapped. This result is in agreement with those by Grishin et al. (2024). Once the thermal torque from Eq. (34) is added to the updated migration torque of Eq. (30), the bottom panel in Fig. 8 shows that we again obtain migration traps. In the SG03 AGN disc, we find two migration traps for a $M=10^{6}\,{\rm M}_{\odot}$ central BH and a $10M_{\odot}$ migrator occurring at $r\approx 1.4\times 10^{3}R_{\rm s}=1.4\times 10^{-4}$ pc and $r\approx 6.8\times 10^{4}R_{\rm s}=6.5\times 10^{-3}$ pc.
When considering the Thompson et al. (2005) model, we obtain a larger number of migration traps, irrespective of the torque prescriptions adopted and the inclusion of the thermal torque contribution to the net torque. For the migration torque by Jiménez & Masset (2017) (the top panel in Fig. 8), we find migration traps form in the TQM05 disc when the gradient $\mathrm{d}\ln\Sigma_{\rm g}/\mathrm{d}\ln r$ discretely changes values, as can be seen in the lower panels of Fig. 8 at $r\approx 2.5\times 10^{3}R_{\rm s}$ and $r\approx 8.3\times 10^{3}R_{\rm s}$. When both migration and thermal torques are considered, we find traps at $r\approx 2.5\times 10^{3}R_{\rm s}$, $r\approx 8.4\times 10^{3}R_{\rm s}$ and $r\approx 2.6\times 10^{6}R_{\rm s}$ for the Thompson et al. (2005) model.
@@ Acknowledgements
We thank Maria Paola Vaccaro, Andrea Derdzinski, Evgeni Grishin, Ari Laor, Hiromichi Tagawa, Shmuel Gilbaum, Sean McGee, Cressida Cleland for discussions. D.Gan. and D.Ger. are supported by ERC Starting Grant No. 945155–GWmining, Cariplo Foundation Grant No. 2021-0555, MUR PRIN Grant No. 2022-Z9X4XS, and the ICSC National Research Centre funded by NextGenerationEU. A.A.T. acknowledges support from JSPS KAKENHI Grant No. 21K13914 and from the European Union’s Horizon 2020 and Horizon Europe research and innovation programs under the Marie Skłodowska-Curie grant agreements No. 847523 and No. 101103134. D.Ger. is supported by MSCA Fellowships No. 101064542–StochRewind and No. 101149270–ProtoBH. Computational work was performed at CINECA with allocations through INFN and Bicocca.